$SPAD/src/input exprode.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool exprode.output
)set message test on
)set message auto off
)clear all
)set streams calculate 7

———-

2 The problem

We will solve y′′′ = sin(y′′) ∗ exp(y) + cos(x) subject to y(0) = 1, y′(0) = 0, y′′(0) = 0 — * —

--S 1 of 13
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 1
--S 2 of 13
eq := differentiate(y x, x, 3) - sin differentiate(y x, x, 2) * exp y x
           = cos x
--R
--R
--R         ,,,        y(x)     ,,
--R   (2)  y   (x) - %e    sin(y  (x))= cos(x)
--R
--R                                          Type: Equation(Expression(Integer))
--E 2
--S 3 of 13
seriesSolve(eq, y, x = 0, [1, 0, 0])
--R
--I   Compiling function %B with type List(UnivariateTaylorSeries(
--I      Expression(Integer),x,0)) -> UnivariateTaylorSeries(Expression(

--I      Integer),x,0)
--R
--R   (3)
--R                          2            3              4      2
--R         1  3   %e  4   %e  - 1  5   %e  - 2%e  6   %e  - 8%e  + 4%e + 1  7
--R     1 + - x  + -- x  + ------- x  + --------- x  + -------------------- x
--R         6      24        120           720                 5040
--R   +
--R        8
--R     O(x )
--R                        Type: UnivariateTaylorSeries(Expression(Integer),x,0)
--E 3

———-

Airy, isn’t it?

— * —

--S 4 of 13
airy := differentiate(y x, x, 2) - x * y x
--R
--R
--R         ,,
--R   (4)  y  (x) - x y(x)
--R
--R                                                    Type: Expression(Integer)
--E 4
--S 5 of 13
seriesSolve(airy, y, x = 0, [a0, a1])
--R
--I   Compiling function %D with type List(UnivariateTaylorSeries(
--I      Expression(Integer),x,0)) -> UnivariateTaylorSeries(Expression(
--I      Integer),x,0)
--R
--R                    a0  3   a1  4    a0  6    a1  7      8
--R   (5)  a0 + a1 x + -- x  + -- x  + --- x  + --- x  + O(x )
--R                     6      12      180      504
--R                        Type: UnivariateTaylorSeries(Expression(Integer),x,0)
--E 5

———-

We can solve around other points than x = 0 — * —

--S 6 of 13
seriesSolve(airy, y, x = 1, [a0, a1])
--R
--I   Compiling function %F with type List(UnivariateTaylorSeries(
--I      Expression(Integer),x,1)) -> UnivariateTaylorSeries(Expression(

--I      Integer),x,1)
--R
--R   (6)
--R                      a0        2   a1 + a0        3   2a1 + a0        4
--R     a0 + a1(x - 1) + -- (x - 1)  + ------- (x - 1)  + -------- (x - 1)
--R                       2               6                  24
--R   +
--R     a1 + 4a0        5   6a1 + 5a0        6   11a1 + 9a0        7            8
--R     -------- (x - 1)  + --------- (x - 1)  + ---------- (x - 1)  + O((x - 1) )
--R        120                 720                  5040
--R                        Type: UnivariateTaylorSeries(Expression(Integer),x,1)
--E 6

———-

System of equations for tan(t) and sec(t)

— * —

--S 7 of 13
x := operator 'x
--R
--I   Compiled code for %F has been cleared.
--I   Compiled code for %D has been cleared.
--I   Compiled code for %B has been cleared.
--R
--R   (7)  x
--R                                                          Type: BasicOperator
--E 7
--S 8 of 13
eq1 := differentiate(x t, t) = 1 + x(t)**2
--R
--R
--R         ,         2
--R   (8)  x (t)= x(t)  + 1
--R
--R                                          Type: Equation(Expression(Integer))
--E 8
--S 9 of 13
eq2 := differentiate(y t, t) = x(t) * y(t)
--R
--R
--R         ,
--R   (9)  y (t)= x(t)y(t)
--R
--R                                          Type: Equation(Expression(Integer))
--E 9
--S 10 of 13

seriesSolve([eq2, eq1], [x, y], t = 0, [y 0 = 1, x 0 = 0])
--R
--I   Compiling function %H with type List(UnivariateTaylorSeries(
--I      Expression(Integer),t,0)) -> UnivariateTaylorSeries(Expression(
--I      Integer),t,0)
--I   Compiling function %I with type List(UnivariateTaylorSeries(
--I      Expression(Integer),t,0)) -> UnivariateTaylorSeries(Expression(
--I      Integer),t,0)
--R
--R              1  3    2  5    17  7      8      1  2    5  4    61  6      8
--R   (10)  [t + - t  + -- t  + --- t  + O(t ),1 + - t  + -- t  + --- t  + O(t )]
--R              3      15      315                2      24      720
--R                  Type: List(UnivariateTaylorSeries(Expression(Integer),t,0))
--E 10

———-

System of equations for a damped pendulum of mass and length 1:

— * —

--S 11 of 13
eq1 := differentiate(x t, t) = y t
--R
--R
--R          ,
--R   (11)  x (t)= y(t)
--R
--R                                          Type: Equation(Expression(Integer))
--E 11
--S 12 of 13
eq2 := differentiate(y t, t) = - g * sin(x t) - c * y t
--R
--R
--R          ,
--R   (12)  y (t)= - g sin(x(t)) - c y(t)
--R
--R                                          Type: Equation(Expression(Integer))
--E 12
--S 13 of 13
seriesSolve([eq1, eq2], [x, y], t = 0, [y 0 = a, x 0 = b])
--R
--I   Compiling function %K with type List(UnivariateTaylorSeries(
--I      Expression(Integer),t,0)) -> UnivariateTaylorSeries(Expression(
--I      Integer),t,0)
--I   Compiling function %L with type List(UnivariateTaylorSeries(
--I      Expression(Integer),t,0)) -> UnivariateTaylorSeries(Expression(
--I      Integer),t,0)
--R

--R   (13)
--R   [
--R                                                                    2
--R                 - g sin(b) - a c  2   c g sin(b) - a g cos(b) + a c   3
--R       b + a t + ---------------- t  + ------------------------------ t
--R                         2                            6
--R     +
--R         2             2    2                               3
--R       (g cos(b) + (- c  + a )g)sin(b) + 2a c g cos(b) - a c   4
--R       ------------------------------------------------------ t
--R                                 24
--R     +
--R                   2      2          2           3     2                2      2
--R             - 3a g sin(b)  + (- 2c g cos(b) + (c  - 4a c)g)sin(b) + a g cos(b)
--R           +
--R                    2    3               4
--R             (- 3a c  + a )g cos(b) + a c
--R        /
--R           120
--R      *
--R          5
--R         t
--R     +
--R               3      3          2      2
--R             3g sin(b)  + 13a c g sin(b)
--R           +
--R                 3      2      2      2  2             4      2 2    4
--R             (- g cos(b)  + (3c  - 11a )g cos(b) + (- c  + 11a c  - a )g)sin(b)
--R           +
--R                     2      2        3     3                5
--R             - 3a c g cos(b)  + (4a c  - 7a c)g cos(b) - a c
--R        /
--R           720
--R      *
--R          6
--R         t
--R     +
--R                    3      3         3                 2      3  2       2
--R             - 13c g sin(b)  + (33a g cos(b) + (- 38a c  + 15a )g )sin(b)
--R           +
--R                    3      2        3      2   2           5      2 3      4
--R               (3c g cos(b)  + (- 4c  + 75a c)g cos(b) + (c  - 26a c  + 11a c)g)
--R            *
--R               sin(b)
--R           +
--R                  3      3        2      3  2      2
--R             - a g cos(b)  + (6a c  - 11a )g cos(b)
--R           +
--R                    4      3 2    5               6
--R             (- 5a c  + 32a c  - a )g cos(b) + a c

--R        /
--R           5040
--R      *
--R          7
--R         t
--R     +
--R          8
--R       O(t )
--R     ,
--R
--R                                                              2
--R                                 c g sin(b) - a g cos(b) + a c   2
--R       a + (- g sin(b) - a c)t + ------------------------------ t
--R                                                2
--R     +
--R         2             2    2                               3
--R       (g cos(b) + (- c  + a )g)sin(b) + 2a c g cos(b) - a c   3
--R       ------------------------------------------------------ t
--R                                  6
--R     +
--R                   2      2          2           3     2                2      2
--R             - 3a g sin(b)  + (- 2c g cos(b) + (c  - 4a c)g)sin(b) + a g cos(b)
--R           +
--R                    2    3               4
--R             (- 3a c  + a )g cos(b) + a c
--R        /
--R           24
--R      *
--R          4
--R         t
--R     +
--R               3      3          2      2
--R             3g sin(b)  + 13a c g sin(b)
--R           +
--R                 3      2      2      2  2             4      2 2    4
--R             (- g cos(b)  + (3c  - 11a )g cos(b) + (- c  + 11a c  - a )g)sin(b)
--R           +
--R                     2      2        3     3                5
--R             - 3a c g cos(b)  + (4a c  - 7a c)g cos(b) - a c
--R        /
--R           120
--R      *
--R          5
--R         t
--R     +
--R                    3      3         3                 2      3  2       2
--R             - 13c g sin(b)  + (33a g cos(b) + (- 38a c  + 15a )g )sin(b)
--R           +
--R                    3      2        3      2   2           5      2 3      4
--R               (3c g cos(b)  + (- 4c  + 75a c)g cos(b) + (c  - 26a c  + 11a c)g)

--R            *
--R               sin(b)
--R           +
--R                  3      3        2      3  2      2
--R             - a g cos(b)  + (6a c  - 11a )g cos(b)
--R           +
--R                    4      3 2    5               6
--R             (- 5a c  + 32a c  - a )g cos(b) + a c
--R        /
--R           720
--R      *
--R          6
--R         t
--R     +
--R                   4             2      2  3       3
--R             (- 33g cos(b) + (38c  - 78a )g )sin(b)
--R           +
--R                        3               3       3   2       2
--R             (- 228a c g cos(b) + (94a c  - 164a c)g )sin(b)
--R           +
--R                  4      3        2       2  3      2
--R                 g cos(b)  + (- 6c  + 102a )g cos(b)
--R               +
--R                  4       2 2      4  2             6      2 4      4 2    6
--R               (5c  - 334a c  + 57a )g cos(b) + (- c  + 57a c  - 76a c  + a )g
--R            *
--R               sin(b)
--R           +
--R                   3      3           3       3   2      2
--R             4a c g cos(b)  + (- 10a c  + 108a c)g cos(b)
--R           +
--R                  5       3 3      5                7
--R             (6a c  - 122a c  + 16a c)g cos(b) - a c
--R        /
--R           5040
--R      *
--R          7
--R         t
--R     +
--R          8
--R       O(t )
--R     ]
--R                  Type: List(UnivariateTaylorSeries(Expression(Integer),t,0))
--E 13
)spool
)lisp (bye)

———-

References^{[1] nothing}

10